OpenStudy (anonymous):

simlify the ratio: 1.16 to 3.16 type as a fraction

OpenStudy (anonymous):

1 4/25 : 3 4/25 29/25 : 79/25 29/79

OpenStudy (anonymous):

now 432 to 153

OpenStudy (anonymous):

But I haven't got a medal for the last two :-(

OpenStudy (anonymous):

get 1 from my side buddy @ few

OpenStudy (anonymous):

now 432 to 153

OpenStudy (anonymous):

both can be divided by 3

OpenStudy (anonymous):

\[\frac{432}{153} = \frac{144}{51} = \frac{48}{17} = 2\frac{14}{17}\]

OpenStudy (anonymous):

no i need it as a ratio fraction

OpenStudy (anonymous):

so take 2nd last one 48 : 17

OpenStudy (anonymous):

is it ok !

OpenStudy (anonymous):

Hahahaha thanks sheg! :-P